15. An unstable high-energy particle enters a detector and leaves a track 1.13 mm long before it decays. Its speed relative to the detector was 0.882c. What is its proper lifetime in seconds? That is, how long would the particle have lasted before decay had it been at rest with respect to the detector?
Answer #1

Answer: to 1/3x10?m 0.8820 tactual at =+/1-60.882) 2014715 tactual = 0.471 X 1.13x103 0.882 * 3x108 li tactual = 2.01x10 12 s